OpenStudy (anonymous):

what is the answer to the following equation? y=-x+7 and y=1/2x+1

OpenStudy (anonymous):

You can solve by elimination, graphing or substitution for this system of equations. In this case, I think that it would be easiest to solve by substituion. Simply enough, substituting is isolating one of the variables of one of the equations to one side and plugging it into the 2nd equation for that variable so that there is only one variable in the 2nd equation. Here, I will substitute y = -x + 7 into y = 1/2x + 1. -x + 7 = .5x + 1 Then all you have to do is isolate the x variable to one side and the constants to the other. -0.5x = -6 x = 12 Sub x = 12 into either the 1st or second equation and solve :) y = -(12) + 7 y = -5 Therefore, the answer in an ordered pair is (12, -5) Double check by plugging the values of x and y into both of the equations.